Solving by Substitution
{x+2y=14
y=3x-14

Answer by solver15(27)

{x+2y=14
y=3x-14
plug in the y=3x-14 to x+2y=14
x+2(3x-14)=14
x+6x-28=14
and then combine like terms and add 28 to both sides of the equation
7x-28=14
+28 +28
7x/7=42/7
x=6
piug in 6 back in to the previous equation
y=3(6)-14
y=18-14
y=4
oyur set of solution is (6,4)